OpenStudy (alex136):

Write the expression (3 − 9i) + (4 + 5i) as a complex number in standard form.?

OpenStudy (alex136):

I think the answer is 7-4i. Am I correct?

OpenStudy (wendy.ivette11714):

You can just remove the parenthesis and simplify by gathering like terms and you go from 3 - 9i + 4 + 5i = 3 + 4 + 5i - 9i (commutative property of addition) = 7 - 4i Yes you are correct :)

OpenStudy (alex136):

Thank you :)

OpenStudy (wendy.ivette11714):

Your Welcome :)